arricakey:

2(3b - 2) = 2(2b + 12)

snowflake0531:

distributive property \[ a(b+c) = ab+ac \\ a(b-c) = ab-ac\]

snowflake0531:

\[2(3b-2) = 2(2b+12)\\6b-4=4b+24\\2b=28\]